# Stationary and Non-Stationary Transition Probabilities in Decision Making: Modeling COVID-19 Dynamics

### v1

@@ Abstract
This study explores the complexities of stationary and non-stationary transition probabilities within the framework of a Markov Decision Process (MDP), specifically applied to COVID-19. The research highlights the critical role these probabilities play in accurately modeling disease dynamics and informing evidence-based decisions by policymakers and public health authorities. By incorporating both stationary transition probabilities (which assume constant rates of state changes) and non-stationary transition probabilities (which adapt to evolving conditions), the findings are pivotal for offering practical insights into optimizing resource allocation and intervention strategies to mitigate the pandemic’s impacts. The structured analysis within this paper includes a detailed model description, derivation of balanced systems, and formulation of transition probabilities, all contextualized within a COVID-19 scenario. These contributions are invaluable for enhancing the efficacy of pandemic response strategies, ultimately improving public health outcomes and economic efficiency.
@@ 1 Introduction
Infectious disease modeling has a rich history, evolving from simple mathematical frameworks to sophisticated models incorporating various biological and social factors. To provide solid estimates, models need to be properly calibrated based on empirical evidence (see for instance [5, 8, 9]). One of the earliest models, the Reed-Frost model, employed a discrete-time approach to describe the probability of an individual being infected in successive time intervals. This model laid the foundation for subsequent studies that sought to capture the complexity of epidemic spread through probabilistic means. For some applications and generalization of the Reed-Frost model, see for instance [3, 14, 17, 19, 21, 24].
@@
$\mathbb{P}_{i,j}(t)=\mathbb{P}(X_{q+\triangle q}=j|X_{q}=i,t)$
@@
Stationary models serve as a useful baseline for evaluating the effects of new interventions by providing a consistent reference point. However, these models may misestimate impacts if the effects of interventions or the underlying disease dynamics change over time. Policymakers can use stationary models to identify deviations from expected outcomes, which may signal a need for policy adjustments. In contrast, non-stationary models offer a more dynamic approach by incorporating time-varying transition probabilities. This adaptability allows for real-time adjustments based on the latest data and evolving conditions. By reflecting changes in disease patterns, intervention effectiveness, and external factors, we anticipate that non-stationary models will enable more responsive and accurate decision-making.
This paper is meticulously structured to offer a comprehensive understanding of our findings. Here is an overview of its organization: Section 2, Preliminaries, lays the groundwork by providing essential definitions and theoretical underpinnings necessary for this study. Section 3 forms the core of the paper and is divided into four subsections. Subsection 3.1 describes the model in detail. Subsection 3.2 derives the balanced system associated to our model. In Subsection 3.3, we formulate stationary transition probabilities, while Subsection 3.4 addresses the formulation of non-stationary transition probabilities. Each of these subsections explores the specifics of how these probabilities are computed and their implications for disease modeling, particularly in the context of a COVID-19 Scenario. In Section 4, we present a simulation of COVID-19 spread over a 365-day period, providing a thorough representation of the disease’s distribution and stabilization within the population. Finally, the paper concludes with some remarks in Section 5.
@@ 2.1 Chain-binomial Models
  • $S_{t}+\displaystyle\sum_{\tau=0}^{t}I_{\tau}=N,~{}~{}t=0,1,\cdots$: at any time $t$, the the number of susceptible plus the number of infectious since time 00 gives the total population.
@@ 3.1 Model Description
$\left\{\begin{array}[]{rlllll}\displaystyle\frac{dS}{dt}&=&-\left(\zeta_{ia,s}I_{a}+\zeta_{is,s}I_{s}+\zeta_{h,s}H\right)\displaystyle\frac{S}{N}+\tau_{r,s}R\\ {}\hfil\\ \displaystyle\frac{dE}{dt}&=&\left(\zeta_{ia,s}I_{a}+\zeta_{is,s}I_{s}+\zeta_{h,s}H\right)\displaystyle\frac{S}{N}-\tau_{e,ia}E\\ {}\hfil\\ \displaystyle\frac{dI_{a}}{dt}&=&\tau_{e,ia}E-(\tau_{ia,is}+\tau_{ia,r})I_{a}\\ {}\hfil\\ \displaystyle\frac{dI_{s}}{dt}&=&\tau_{ia,is}I_{a}-(\tau_{is,r}+\tau_{is,h}+\tau_{is,d})I_{s}\\ {}\hfil\\ \displaystyle\frac{dH}{dt}&=&\tau_{is,h}I_{s}-(\tau_{h,r}+\tau_{h,d})H\\ {}\hfil\\ \displaystyle\frac{dR}{dt}&=&\tau_{is,r}I_{s}+\tau_{ia,r}I_{a}+\tau_{h,r}H-\tau_{r,s}R\\ {}\hfil\\ \displaystyle\frac{dD}{dt}&=&\tau_{h,d}H+\tau_{is,d}I_{s}\end{array}\right.,$ (3.1)
@@ 3.3 Stationary Transition Probabilities
    $\begin{array}[]{rlllll}\mathbb{P}(\overline{c}_{k})&=&\mathbb{P}(\overline{m}_{k}\vee(m_{k}\wedge\overline{g}_{k}))\\ {}\hfil\\ &=&\mathbb{P}(\overline{m}_{k})+\mathbb{P}(m_{k}\wedge\overline{g}_{k})\\ {}\hfil\\ &=&\mathbb{P}(\overline{m}_{k})+\mathbb{P}(\overline{g}_{k}|m_{k})\mathbb{P}(m_{k}).\end{array}$
@@
    $\begin{array}[]{rlllll}\mathbb{P}(\overline{c}_{k})&=&\mathbb{P}(\overline{m}_{k})+(1-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s}))\mathbb{P}(m_{k})\\ {}\hfil\\ &=&\displaystyle\frac{|S|_{k}+|R|_{k}}{N-|D|_{k}}+(1-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s}))\displaystyle\frac{|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ {}\hfil\\ &=&\displaystyle\frac{|S|_{k}+|R|_{k}+|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ {}\hfil\\ &&-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})\displaystyle\frac{|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ {}\hfil\\ &=&1-\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}.\end{array}$
@@
The transition probabilities from a state to another are given by:
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S)&=&\begin{pmatrix}|S|_{k}\\ ~{}\\ \delta_{1}\end{pmatrix}\left(\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{\delta_{1}}\\ {}\hfil\\ &&\left(1-\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{|S|_{k}-\delta_{1}}\end{array}$ (3.14)
@@
    $\mathbb{P}(I_{s}|I_{a})=\displaystyle\frac{\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{3}\end{pmatrix}\tau_{ia,is}^{\delta_{3}}(1-\tau_{ia,is})^{\delta_{4}}}{{\mathcal{M}}_{|I_{a}|_{k},\delta_{3},\delta_{4}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-(\tau_{ia,is}+\tau_{ia,r}))^{|I_{a}|_{k}-(\delta_{3}+\delta_{4})}}.$ (3.16)
@@
    $\mathbb{P}(R|I_{a})=\displaystyle\frac{\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{4}\end{pmatrix}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,r})^{\delta_{3}}}{{\mathcal{M}}_{|I_{a}|_{k},\delta_{3},\delta_{4}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-(\tau_{ia,is}+\tau_{ia,r}))^{|I_{a}|_{k}-(\delta_{3}+\delta_{4})}}.$ (3.17)
@@
    $\mathbb{P}(R|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{5}\end{pmatrix}\tau_{is,r}^{\delta_{5}}(1-\tau_{is,r})^{\delta_{6}+\delta_{7}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}(1-(\tau_{is,r}+\tau_{is,h}+\tau_{is,d}))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}}.$ (3.18)
@@
    $\mathbb{P}(H|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{6}\end{pmatrix}\tau_{is,h}^{\delta_{6}}(1-\tau_{is,h})^{\delta_{5}+\delta_{7}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}(1-(\tau_{is,r}+\tau_{is,h}+\tau_{is,d}))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}}.$ (3.19)
@@
    $\mathbb{P}(D|I_{s})=\displaystyle\frac{\begin{pmatrix}\delta_{5}+\delta_{6}+\delta_{7}\\ ~{}\\ \delta_{7}\end{pmatrix}\tau_{is,d}^{\delta_{7}}(1-\tau_{is,d})^{\delta_{5}+\delta_{6}}}{{\mathcal{M}}_{|I_{s}|_{k},\delta_{5},\delta_{6},\delta_{7}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}(1-(\tau_{is,r}+\tau_{is,h}+\tau_{is,d}))^{|I_{s}|_{k}-(\delta_{5}+\delta_{6}+\delta_{7})}}.$ (3.20)
@@
    $\mathbb{P}(D|H)=\displaystyle\frac{\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{8}\end{pmatrix}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,d})^{\delta_{9}}}{{\mathcal{M}}_{|H|_{k},\delta_{8},\delta_{9}}\tau_{h,d}^{\delta_{8}}\tau_{h,r}^{\delta_{9}}(1-(\tau_{h,d}+\tau_{h,r}))^{|H|_{k}-(\delta_{8}+\delta_{9})}}.$ (3.21)
@@
    $\mathbb{P}(R|H)=\displaystyle\frac{\begin{pmatrix}\delta_{8}+\delta_{9}\\ ~{}\\ \delta_{9}\end{pmatrix}\tau_{h,r}^{\delta_{9}}(1-\tau_{h,r})^{\delta_{8}}}{{\mathcal{M}}_{|H|_{k},\delta_{8},\delta_{9}}\tau_{h,d}^{\delta_{8}}\tau_{h,r}^{\delta_{9}}(1-(\tau_{h,d}+\tau_{h,r}))^{|H|_{k}-(\delta_{8}+\delta_{9})}}.$ (3.22)
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S)&=&\mathbb{P}({\mathcal{V}}_{\delta_{1}})\\ {}\hfil\\ &=&\mathcal{B}(\delta_{1};|S|_{k},\mathbb{P}(g_{k}))\\ {}\hfil\\ &=&\begin{pmatrix}|S|_{k}\\ ~{}\\ \delta_{1}\end{pmatrix}(\mathbb{P}(g_{k}))^{\delta_{1}}(1-\mathbb{P}(g_{k}))^{|S|_{k}-\delta_{1}}.\end{array}$
@@
    $\begin{array}[]{rllll}\mathbb{P}(I_{a}|E)&=&\mathbb{P}({\mathcal{V}}_{\delta_{2}})\\ {}\hfil\\ &=&\mathcal{B}(\delta_{2};|E|_{k},\tau_{e,ia}).\end{array}$
@@
    where $\mathbb{P}(I_{a})=\mathbb{P}({\mathcal{V}}_{\delta_{3},\delta_{4}})$ and $\mathbb{P}(I_{s}\cap I_{a})=\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{3}\end{pmatrix}\tau_{ia,is}^{\delta_{3}}(1-\tau_{ia,is})^{\delta_{4}}$.
@@
    where $\mathbb{P}(I_{a})=\mathbb{P}({\mathcal{V}}_{\delta_{3},\delta_{4}})$ and $\mathbb{P}(R\cap I_{a})=\begin{pmatrix}\delta_{3}+\delta_{4}\\ ~{}\\ \delta_{4}\end{pmatrix}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,r})^{\delta_{3}}$.

### v2

@@ Abstract
This study introduces a comparative modeling framework using stationary and non-stationary transition probabilities within a Markov Decision Process (MDP) to assess COVID-19 disease dynamics. Stationary transition probabilities assume constant transition rates, while non-stationary transitions reflect time-dependent behaviors including policy interventions or behavioral changes. We develop a comprehensive compartmental model with transitions based on binomial and multinomial processes. Mathematical models for both stationary and non-stationary transition frameworks are developed and simulated over a 365-day period to emphasize dynamic variations in epidemic outcomes. Our findings highlight the significance of non-stationary modeling in accurately representing the dynamic characteristics of pandemic situations and provide recommendations for optimizing public health interventions under uncertainty. This comparative analysis offers useful information for epidemiological modeling and decision-making in dynamic risk environments.
@@ 1 Introduction
Infectious disease modeling has a rich history, evolving from simple mathematical frameworks to sophisticated models incorporating various biological and social factors. To provide solid estimates, models need to be properly calibrated based on empirical evidence (see for instance [5, 8, 9]). One of the earliest models, the Reed-Frost model, employed a discrete-time approach to describe the probability of an individual being infected in successive time intervals. This model laid the foundation for subsequent studies that sought to capture the complexity of epidemic spread through probabilistic means. For some applications and generalization of the Reed-Frost model, see for instance [3, 14, 17, 19, 21, 24]. Classical compartmental models such as SIR and SEIR generally assume constant transition rates between disease states [2, 12]. However, for pandemics such as Covid-19, these assumptions often break down due to dynamic changes such as: behavioral changes, fluctuation in healthcare capacities or public health policies [5, 11, 20, 25]. Consequently, incorporating non-stationary transition probabilities enables a more realistic epidemic dynamics.
@@
$\mathbb{P}_{i,j}(t)=\mathbb{P}(X_{q+\triangle q}=j|X_{q}=i,~t)$
@@
Stationary models serve as a useful baseline for evaluating the effects of new interventions by providing a consistent reference point. However, these models may misestimate impacts if the effects of interventions or the underlying disease dynamics change over time. Policymakers can use stationary models to identify deviations from expected outcomes, which may signal a need for policy adjustments. In contrast, non-stationary models offer a more dynamic approach by incorporating time-varying transition probabilities. This adaptability allows for real-time adjustments based on the latest data and evolving conditions. By reflecting changes in disease patterns, intervention effectiveness, and external factors, we anticipate that non-stationary models will enable more responsive and accurate decision-making. Our primary goal is to assess how different assumptions about transition dynamics impact model outcomes and their relevance to public health policy.
The main contributions of this study include developing an epidemiological model with seven epidemiological compartments reflecting key clinical stages of COVID-19. We then derive probabilistic transition rules between states using binomial and multinomial distributions, and explicitly distinguish between stationary and non-stationary regimes. We define time-dependent transition parameters using logistic and quadratic forms to reflect real-world dynamics such as behavioral adaptation or resource saturation. Finally, we simulate COVID-19 dynamics under both transition assumptions and compare outcomes over a 365-day period, illustrating the impact of model choice on disease spread over time. We offer policy-relevant interpretations of the modeling differences, particularly with respect to intervention timing and uncertainty in parameter estimation.
This paper is meticulously structured to offer a comprehensive understanding of our findings. Here is an overview of its organization: Section 2 presents key probabilistic tools and foundational concepts. Section 3 details the compartmental model and develops transition probability expressions for both stationary and non-stationary settings. Section 4 presents simulations comparing the two regimes. Section 5 concludes with a discussion of practical implications and directions for future work.
@@ 2.1 Chain-binomial Models
  • $S_{t}+\displaystyle\sum_{\tau=0}^{t}I_{\tau}=N,~~t=0,1,\cdots$: at any time $t$, the the number of susceptible plus the number of infectious since time 0 gives the total population.
@@ 3.1 Model Description
$\left\{\begin{array}[]{rlllll}\displaystyle\frac{dS}{dt}&=&-\left(\zeta_{ia,s}I_{a}+\zeta_{is,s}I_{s}+\zeta_{h,s}H\right)\displaystyle\frac{S}{N}+\tau_{r,s}R\\ \\ \displaystyle\frac{dE}{dt}&=&\left(\zeta_{ia,s}I_{a}+\zeta_{is,s}I_{s}+\zeta_{h,s}H\right)\displaystyle\frac{S}{N}-\tau_{e,ia}E\\ \\ \displaystyle\frac{dI_{a}}{dt}&=&\tau_{e,ia}E-(\tau_{ia,is}+\tau_{ia,r})I_{a}\\ \\ \displaystyle\frac{dI_{s}}{dt}&=&\tau_{ia,is}I_{a}-(\tau_{is,r}+\tau_{is,h}+\tau_{is,d})I_{s}\\ \\ \displaystyle\frac{dH}{dt}&=&\tau_{is,h}I_{s}-(\tau_{h,r}+\tau_{h,d})H\\ \\ \displaystyle\frac{dR}{dt}&=&\tau_{is,r}I_{s}+\tau_{ia,r}I_{a}+\tau_{h,r}H-\tau_{r,s}R\\ \\ \displaystyle\frac{dD}{dt}&=&\tau_{h,d}H+\tau_{is,d}I_{s}\end{array}\right.,$ (3.1)
@@ 3.3 Stationary Transition Probabilities
    $\begin{array}[]{rlllll}\mathbb{P}(\overline{c}_{k})&=&\mathbb{P}(\overline{m}_{k}\vee(m_{k}\wedge\overline{g}_{k}))\\ \\ &=&\mathbb{P}(\overline{m}_{k})+\mathbb{P}(m_{k}\wedge\overline{g}_{k})\\ \\ &=&\mathbb{P}(\overline{m}_{k})+\mathbb{P}(\overline{g}_{k}|m_{k})\mathbb{P}(m_{k}).\end{array}$
@@
    $\begin{array}[]{rlllll}\mathbb{P}(\overline{c}_{k})&=&\mathbb{P}(\overline{m}_{k})+(1-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s}))\mathbb{P}(m_{k})\\ \\ &=&\displaystyle\frac{|S|_{k}+|R|_{k}}{N-|D|_{k}}+(1-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s}))\displaystyle\frac{|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ \\ &=&\displaystyle\frac{|S|_{k}+|R|_{k}+|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ \\ &&-(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})\displaystyle\frac{|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k}}{N-|D|_{k}}\\ \\ &=&1-\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}.\end{array}$
@@
The transition probabilities from a state to another are given by the following, where
$\beta=1-\tau_{is,r}-\tau_{is,h}-\tau_{is,d},~~Q_{1}=|I_{a}|_{k}-\delta_{3}-\delta_{4},~\text{and}~Q_{2}=|I_{s}|_{k}-\delta_{5}-\delta_{6}-\delta_{7}.$
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S)&=&\begin{pmatrix}|S|_{k}\\ ~\\ \delta_{1}\end{pmatrix}\left(\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{\delta_{1}}\\ \\ &&\left(1-\displaystyle\frac{(\zeta_{ia,s}+\zeta_{is,s}+\zeta_{h,s})(|E|_{k}+|I_{a}|_{k}+|I_{s}|_{k}+|H|_{k})}{N-|D|_{k}}\right)^{|S|_{k}-\delta_{1}}\end{array}$ (3.14)
@@
    $\mathbb{P}(I_{s}\mid I_{a})=\frac{\displaystyle\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}}{\displaystyle\sum_{\delta_{3}=0}^{|I_{a}|_{k}}\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}}.$ (3.16)
@@
    $\mathbb{P}(R\mid I_{a})=\frac{\displaystyle\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}}{\displaystyle\sum_{\delta_{4}=0}^{|I_{a}|_{k}}\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}}.$ (3.17)
@@
    $\mathbb{P}(R|I_{s})=\frac{\displaystyle\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}.$ (3.18)
@@
    $\mathbb{P}(H|I_{s})=\frac{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{6}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}{\displaystyle\sum_{\delta_{6}=0}^{|I_{s}|_{k}}\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{6}}\sum_{\delta_{7}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{6}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}.$ (3.19)
@@
    $\mathbb{P}(D|I_{s})=\frac{\displaystyle\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{7}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{7}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}{\displaystyle\sum_{\delta_{7}=0}^{|I_{s}|_{k}}\sum_{\delta_{5}=0}^{|I_{s}|_{k}-\delta_{7}}\sum_{\delta_{6}=0}^{|I_{s}|_{k}-\delta_{5}-\delta_{7}}\binom{|I_{s}|_{k}}{\delta_{5},\delta_{6},\delta_{7},Q_{2}}\tau_{is,r}^{\delta_{5}}\tau_{is,h}^{\delta_{6}}\tau_{is,d}^{\delta_{7}}\beta^{Q_{2}}}.$ (3.20)
@@
    $\mathbb{P}(D|H)=\frac{\displaystyle\sum_{\delta_{9}=0}^{|H|_{k}-\delta_{8}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}^{\delta_{9}}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,r}-\tau_{h,d})^{|H|_{k}-\delta_{8}-\delta_{9}}}{\displaystyle\sum_{\delta_{8}=0}^{|H|_{k}}\sum_{\delta_{9}=0}^{|H|_{k}-\delta_{8}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}^{\delta_{9}}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,r}-\tau_{h,d})^{|H|_{k}-\delta_{8}-\delta_{9}}}.$ (3.21)
@@
    $\mathbb{P}(R|H)=\frac{\displaystyle\sum_{\delta_{8}=0}^{|H|_{k}-\delta_{9}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}^{\delta_{9}}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,r}-\tau_{h,d})^{|H|_{k}-\delta_{8}-\delta_{9}}}{\displaystyle\sum_{\delta_{9}=0}^{|H|_{k}}\sum_{\delta_{8}=0}^{|H|_{k}-\delta_{9}}\binom{|H|_{k}}{\delta_{8},\delta_{9},|H|_{k}-\delta_{8}-\delta_{9}}\tau_{h,r}^{\delta_{9}}\tau_{h,d}^{\delta_{8}}(1-\tau_{h,r}-\tau_{h,d})^{|H|_{k}-\delta_{8}-\delta_{9}}}.$ (3.22)
@@
    $\begin{array}[]{rllll}\mathbb{P}(E|S)&=&\mathbb{P}({\mathcal{V}}_{\delta_{1}})\\ \\ &=&\mathcal{B}(\delta_{1};|S|_{k},\mathbb{P}(g_{k}))\\ \\ &=&\begin{pmatrix}|S|_{k}\\ ~\\ \delta_{1}\end{pmatrix}(\mathbb{P}(g_{k}))^{\delta_{1}}(1-\mathbb{P}(g_{k}))^{|S|_{k}-\delta_{1}}.\end{array}$
@@
    $\begin{array}[]{rllll}\mathbb{P}(I_{a}|E)&=&\mathbb{P}({\mathcal{V}}_{\delta_{2}})\\ \\ &=&\mathcal{B}(\delta_{2};|E|_{k},\tau_{e,ia}).\end{array}$
@@
    where
    $\mathbb{P}(I_{s}\cap I_{a})=\displaystyle\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}$ (3.24)
    is obtained by summing over all possible values of $\delta_{4}$, and
    $\mathbb{P}(I_{a})=\sum_{\delta_{3}=0}^{|I_{a}|_{k}}\sum_{\delta_{4}=0}^{|I_{a}|_{k}-\delta_{3}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}.$ (3.25)
@@
    where
    $\mathbb{P}(R\cap I_{a})=\displaystyle\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}$ (3.26)
    is obtained by summing over all possible values of $\delta_{3}$, and
    $\mathbb{P}(I_{a})=\sum_{\delta_{4}=0}^{|I_{a}|_{k}}\sum_{\delta_{3}=0}^{|I_{a}|_{k}-\delta_{4}}\binom{|I_{a}|_{k}}{\delta_{3},\delta_{4},Q_{1}}\tau_{ia,is}^{\delta_{3}}\tau_{ia,r}^{\delta_{4}}(1-\tau_{ia,is}-\tau_{ia,r})^{Q_{1}}.$ (3.27)